# Nonstandard Universes

### v1

@@ 6.4.1 Construction of Polysaturated Extensions via Good Ultrapowers
Recall the ultrapower construction of a nonstandard extension $V(X)\overset{*}{\hookrightarrow}V(^{*}X)$: We start with a base set $X$ and an ultrafilter $\mathcal{U}$ over a set $I$, where $I$ is chosen so that the ultrapower $Y:=X^{I}/\mathcal{U}$ is another base set. We define relations $=_{\mathcal{U}},\in_{\mathcal{U}}$ on $V(X)^{I}$ by
@@
[…]
@@ 6.5.1 Limits of Chains of Superstructures
Now define a sequence of sets $(W_{n})_{n\in\mathbb{N}}$ by induction, as follows: First, if $(a,\alpha)\in P_{0}$, let $[a,\alpha]:=\{(b,\beta):(b,\beta)\sim(a,\alpha)\}$. By Lemma 6.23, $[a,\alpha]\subseteq P_{0}$. Then let
@@
Now assume that $[b,\beta]$ and $W_{k}$ have already been defined for all $k<n$ and $(b,\beta)\in P_{k}$. Now define $[a,\alpha]$ for $(a,\alpha)\in P_{n}$, and define $W_{n}$ by
$[a,\alpha]:=\{[b,\beta]:(b,\beta)E(a,\alpha)\},\qquad W_{n}:=\{[a,\alpha]:(a,\alpha)\in P_{n}\}.$
(Observe that if $(b,\beta)E(a,\alpha)$, then $(b,\beta)\in P_{n-1}$ by Lemma 6.23, so that $[b,\beta]$ has already been defined.) Then define
@@
The elements of $W_{0}=X^{\lambda}$ act as atoms. The elements of $W-W_{0}$ act as sets. Observe that $W$ is transitive over sets: If $x\in[a,\alpha]$, where $[a,\alpha]\in W-W_{0}$ is a set, then $x=[b,\beta]$ for some $(b,\beta)E(a,\alpha)$ and hence $x\in W$ also. Note also that $W_{n}\subseteq V_{n}(X^{\lambda})$, for all $n\in\mathbb{N}$.
@@
    If $\alpha\leq\beta<\lambda$ and $a\in V(X^{\alpha})$, then $[a,\alpha]=[\iota_{\alpha\beta}(a),\beta]$.
@@
    If $\alpha,\beta<\lambda$, then $[b,\beta]\in[X^{\alpha},\alpha]$ if and only if $b\in X^{\beta}$.
@@
$[\iota_{\alpha\beta}(a),\beta]$ =$=\{(c,\gamma):(c,\gamma)\sim(\iota_{\alpha\beta}(a),\beta)\}$
=$=\{(c,\gamma):\exists\delta\geq\beta,\gamma\;\Big{(}\iota_{\gamma\delta}(c)=\iota_{\beta\delta}(\iota_{\alpha\beta}(a))=\iota_{\alpha\delta}(a)\Big{)}\}$
=$=\{(c,\gamma):(c,\gamma)\sim(a,\alpha)\}$
@@
$[\iota_{\alpha\beta}(a),\beta]$ =$=\{[c,\gamma]:(c,\gamma)E(\iota_{\alpha\beta}(a),\beta)\}$
=$=\{[c,\gamma]:\exists\delta\geq\beta,\gamma\;\Big{(}\iota_{\gamma\delta}(c)\in\iota_{\beta\delta}(\iota_{\alpha\beta}(a))=\iota_{\alpha\delta}(a)\Big{)}\},$
=$=\{[c,\gamma]:(c,\gamma)E(a,\alpha)\},$
@@
=$=\{[b,\beta]:[b,\beta]\in[X^{\alpha},\alpha]\},$
=$=\{[b,\beta]:b\in X^{\beta}=V_{0}(X^{\beta})\},$
=$=\{[b,\beta]:(b,\beta)\in P_{0}\},$
@@
The propositional connectives ,$\land$,$\lnot$ , ¬ are dealt with very easily.
@@
Conversely, suppose that $V(X^{\lambda})\vDash\exists y\in\iota_{\alpha\lambda}(a_{1})\;\psi(y,\iota_{\alpha\lambda}(a_{1}),\dots,\iota_{\alpha\lambda}(a_{n}))$. Then there is $[b,\beta]\in[a_{1},\alpha]$ such that $V(X^{\lambda})\vDash\psi([b,\beta],[a_{1},\alpha],\dots,[a_{n},\alpha])$. Let $\gamma\geq\alpha,\beta$. Then $[b,\beta]=[\iota_{\beta\gamma}(b),\gamma]=\iota_{\gamma\lambda}(\iota_{\beta\gamma}(b))$ and $[a_{i},\alpha]=[\iota_{\alpha\gamma}(a_{i}),\gamma]=\iota_{\gamma\lambda}(\iota_{\alpha\gamma}(a_{i}))$, with $\iota_{\beta\gamma}(b)\in\iota_{\alpha\gamma}(a_{1})$. Thus
@@ 6.5.2 Construction of Polysaturated Extensions via Ultralimits
Recall Theorem 6.7, which states that for every superstructure $V(X)$ there is a set $I$ and an ultrafilter over $I$ such that the induced bounded elementary embedding $*:V(X)\to V(^{*}X)$ is an enlargement, where ${}^{*}X=X^{I}/\mathcal{U}$. The idea behind the proof of the existence of a polysaturated extension is to iterate this construction.
Let $V(X)$ be a superstructure over a base set $X$, and let $\kappa=|V(X)|$ be its cardinality. For $\alpha\leq\beta\leq\kappa^{+}$, we construct superstructures $V(X^{\alpha})$ and bounded elementary embeddings $*_{\alpha\beta}:V(X^{\alpha})\to V(X^{\beta})$ such that
@@
    If $\alpha\leq\kappa^{+}$, then $*_{\alpha\alpha}=\text{id}_{V(X^{\alpha})}$.
@@
    If $\alpha\leq\beta\leq\gamma\leq\kappa^{+}$, then $*_{\beta\gamma}\circ*_{\alpha\beta}=*_{\alpha\gamma}$.
@@
    If $\alpha\leq\beta\leq\kappa^{+}$, then $*_{\alpha\beta}(X^{\alpha})=X^{\beta}$.
@@
We define $V(X^{0})=V(X)$, and $*_{00}=\text{id}_{V(X)}$.
Suppose now that superstructures $V(X^{\alpha})$ and bounded elementary embeddings $*_{\alpha\beta}:V(X^{\alpha})\to V(X^{\beta})$ have already been constructed for $\alpha\leq\beta<\lambda$, such that (i)-(iii) are satisfied. We now consider two cases:
Case 1: $\lambda$ is a limit ordinal. In that case construct $V(X^{\lambda})$ and $*_{\alpha\lambda}$ as a limit of the $V(X^{\alpha})$ and $*_{\alpha\beta}$ for $\alpha,\beta<\lambda$. We have just seen that such a construction yields bounded elementary embeddings which preserve properties (i)-(iii).
Case 2: $\lambda=\gamma+1$ is a successor ordinal. In that case, let $V(X^{\gamma})\overset{*_{\gamma\lambda}}{\to}V(X^{\lambda})$ be an enlargement, as provided by Theorem 6.7. Observe that $X^{\lambda}={}^{*_{\gamma\lambda}}X^{\gamma}$ by construction. For $\alpha<\gamma$, define $*_{\alpha\lambda}=*_{\gamma\lambda}\circ*_{\alpha\gamma}$, and define $*_{\lambda\lambda}=\text{id}_{V(X^{\lambda})}$. It is then straightforward to show that that properties (i)-(iii) are satisfied.
@@
For $(a,\alpha)\in P_{0}$, we define $[a,\alpha]:=\{(b,\beta):(b,\beta)\sim(a,\alpha)\}$ and for $(a,\alpha)\in P-P_{0}$ we put $[a,\alpha]:=\{[b,\beta]:(b,\beta)E(a,\alpha)$. Then we define
@@
$[a,\alpha]=[*_{\alpha\beta}(a),\beta]\quad\text{ for }\alpha\leq\beta<\kappa^{+}.$
@@
$*_{\alpha\kappa^{+}}:V(X^{\alpha})\to V(X^{\kappa^{+}}):a\mapsto[a,\alpha].$
Let $\mathcal{A}\subseteq W_{n}$ be a family of cardinality $\kappa$ with the f.i.p. We will show that $\bigcap\mathcal{A}\neq\varnothing$. Note that each $A\in\mathcal{A}$ is of the form $[a,\alpha]$ for some $a$ in some $V_{n}(X^{\alpha})$. Suppose that
$\mathcal{A}=\{A_{\xi}:\xi<\kappa\}=\{[a_{\xi},\alpha_{\xi}]:\xi<\kappa\}\quad\text{is an enumeration of }\mathcal{A}.$
Let $\beta:=\sup_{\xi<\kappa}\alpha_{\xi}$. Since $\kappa^{+}$ is a regular cardinal and each $\alpha_{\xi}<\kappa^{+}$, we have that $\beta<\kappa^{+}$. Define $a_{\xi}^{\prime}=*_{\alpha_{\xi}\beta}(a_{\xi})\in V_{n}(X^{\beta})$, so that $A_{\xi}=[a_{\xi},\alpha_{\xi}]=[a^{\prime}_{\xi},\beta]$, and let $\mathcal{A}^{\prime}:=\{a^{\prime}_{\xi}:\xi<\kappa\}$ Observe that $*_{\beta\kappa^{+}}(a^{\prime}_{\xi})=[a^{\prime}_{\xi},\beta]=A_{\xi}$. As $\mathcal{A}$ has the f.i.p. and $*_{\beta\kappa^{+}}$ is a bounded elementary embedding, it follows easily that $\mathcal{A}^{\prime}\subseteq V_{n}(X^{\beta})$ has the f.i.p. as well. As $*_{\beta(\beta+1)}:V(X^{\beta})\to V(X^{\beta+1})$ is an enlargement, it follows that
@@
Now if $x\in\bigcap*_{\alpha(\alpha+1}[\mathcal{A}^{\prime}]$, then $x\in*_{\beta(\beta+1)}a_{\xi}^{\prime}$ for all $\xi<\kappa$. It follows that $*_{(\beta+1)\kappa^{+}}(x)\in*_{(\beta+1)\kappa^{+}}\circ*_{\beta(\beta+1)}(a_{\xi}^{\prime})=*_{\beta\kappa^{+}}(a^{\prime}_{\xi})=A_{\xi}$ for all $\xi<\kappa$, and hence that .$*_{(\beta+1)\kappa^{+}}(x)\in\bigcap\mathcal{A}$. Thus $\bigcap\mathcal{A}\neq\varnothing$ whenever $\mathcal{A}\subseteq W_{n}$ is a family of $\leq\kappa$-many sets with the f.i.p.
Observe that $X^{\kappa^{+}}=*_{\alpha\kappa^{+}}(X_{\alpha})$ for all $\alpha<\kappa^{+}$. We now show that each $*_{\alpha\kappa^{+}}:V(X^{\alpha})\to V(X^{\kappa^{+}})$ is $\kappa^{+}$-saturated.
@@
Since $\kappa=|V(X)|$, it follows that if we define $*:=*_{0\kappa^{+}}$ and ${}^{*}X:=X^{\kappa^{+}}$, then the map $*:V(X)\to V({}^{*}X)$ is polysaturated.

### v2

@@ 6.4.1 Construction of Polysaturated Extensions via Good Ultrapowers
Recall the ultrapower construction of a nonstandard extension $V(X)\overset{*}{\hookrightarrow}V(^{*}X)$: We start with a base set $X$ and an ultrafilter $\mathcal{U}$ over a set $I$, where $I$ is chosen so that the ultrapower $Y:=X^{I}/\mathcal{U}$ is another base set. We define relations $_{\mathcal{U}},\in_{\mathcal{U}}$ on $V(X)^{I}$ by
@@
Then the *-map, defined as the composition $V(X)\overset{\iota}{\hookrightarrow}W\overset{\cdot/\mathcal{U}}{\to}V(Y)$, is a transfer map, with $Y={}^{*}X$. If the ultrafilter $\mathcal{U}$ is also countably incomplete, then $V(X)\to V({}^{*}X)$ is a nonstandard framework, in that $\{{}^{*}a:a\in A\}$ is a proper subset of $^{*}a$ whenever $A\in V(X)$ is an infinite set.
Observe that the internal sets are precisely the sets of the form $f/\mathcal{U}$, for $f\in W$. Indeed, if $f\in W$, then $f\in W_{n}$ for some $n\in\mathbb{N}$, and hence $f\in_{\mathcal{U}}c_{V_{n}(X)}$, so that $f/\mathcal{U}\in{}^{*}V_{n}(X)=c_{V_{n}(X)}/\mathcal{U}$, from which it follows that $f/\mathcal{U}$ is internal. Conversely, if $A\in V(Y)$ is internal, then $A\in{}^{*}B=c_{B}/\mathcal{U}$, where $c_{B}/\mathcal{U}:=\{f/\mathcal{U}:f\in W\land f\in_{\mathcal{U}}c_{B}\}$, from which it follows that $A=f/\mathcal{U}$ for some $f\in W$. Thus
@@ 6.5.1 Limits of Chains of Superstructures
Now define a sequence of sets $(W_{n})_{n\in\mathbb{N}}$ by induction, as follows: First, if $(a,\alpha)\in P_{0}$, let $a,\alpha]:=\{(b,\beta):(b,\beta)\sim(a,\alpha)\}$. By Lemma 6.23, $a,\alpha]\subseteq P_{0}$. Then let
@@
Now assume that $b,\beta]$ and $W_{k}$ have already been defined for all $k<n$ and $(b,\beta)\in P_{k}$. Now define $a,\alpha]$ for $(a,\alpha)\in P_{n}$, and define $W_{n}$ by
$a,\alpha]:=\{[b,\beta]:(b,\beta)E(a,\alpha)\},\qquad W_{n}:=\{[a,\alpha]:(a,\alpha)\in P_{n}\}.$
(Observe that if $(b,\beta)E(a,\alpha)$, then $(b,\beta)\in P_{n-1}$ by Lemma 6.23, so that $b,\beta]$ has already been defined.) Then define
@@
The elements of $W_{0}=X^{\lambda}$ act as atoms. The elements of $W-W_{0}$ act as sets. Observe that $W$ is transitive over sets: If $x\in[a,\alpha]$, where $a,\alpha]\in W-W_{0}$ is a set, then $x=[b,\beta]$ for some $(b,\beta)E(a,\alpha)$ and hence $x\in W$ also. Note also that $W_{n}\subseteq V_{n}(X^{\lambda})$, for all $n\in\mathbb{N}$.
@@
    If $\alpha\leq\beta<\lambda$ and $a\in V(X^{\alpha})$, then $a,\alpha]=[\iota_{\alpha\beta}(a),\beta]$.
@@
    If $\alpha,\beta<\lambda$, then $b,\beta]\in[X^{\alpha},\alpha]$ if and only if $b\in X^{\beta}$.
@@
$[\iota_{\alpha\beta}(a),\beta]$ $=\{(c,\gamma):(c,\gamma)\sim(\iota_{\alpha\beta}(a),\beta)\}$
$=\{(c,\gamma):\exists\delta\geq\beta,\gamma\;\Big(\iota_{\gamma\delta}(c)=\iota_{\beta\delta}(\iota_{\alpha\beta}(a))=\iota_{\alpha\delta}(a)\Big)\}$
$=\{(c,\gamma):(c,\gamma)\sim(a,\alpha)\}$
@@
$[\iota_{\alpha\beta}(a),\beta]$ $=\{[c,\gamma]:(c,\gamma)E(\iota_{\alpha\beta}(a),\beta)\}$
$=\{[c,\gamma]:\exists\delta\geq\beta,\gamma\;\Big(\iota_{\gamma\delta}(c)\in\iota_{\beta\delta}(\iota_{\alpha\beta}(a))=\iota_{\alpha\delta}(a)\Big)\},$
$=\{[c,\gamma]:(c,\gamma)E(a,\alpha)\},$
@@
$=\{[b,\beta]:[b,\beta]\in[X^{\alpha},\alpha]\},$
$=\{[b,\beta]:b\in X^{\beta}=V_{0}(X^{\beta})\},$
$=\{[b,\beta]:(b,\beta)\in P_{0}\},$
@@
The propositional connectives $\land,\lnot$ are dealt with very easily.
@@
Conversely, suppose that $V(X^{\lambda})\vDash\exists y\in\iota_{\alpha\lambda}(a_{1})\;\psi(y,\iota_{\alpha\lambda}(a_{1}),\dots,\iota_{\alpha\lambda}(a_{n}))$. Then there is $b,\beta]\in[a_{1},\alpha]$ such that $V(X^{\lambda})\vDash\psi([b,\beta],[a_{1},\alpha],\dots,[a_{n},\alpha])$. Let $\gamma\geq\alpha,\beta$. Then $b,\beta]=[\iota_{\beta\gamma}(b),\gamma]=\iota_{\gamma\lambda}(\iota_{\beta\gamma}(b))$ and $a_{i},\alpha]=[\iota_{\alpha\gamma}(a_{i}),\gamma]=\iota_{\gamma\lambda}(\iota_{\alpha\gamma}(a_{i}))$, with $\iota_{\beta\gamma}(b)\in\iota_{\alpha\gamma}(a_{1})$. Thus
@@ 6.5.2 Construction of Polysaturated Extensions via Ultralimits
Recall Theorem 6.7, which states that for every superstructure $V(X)$ there is a set $I$ and an ultrafilter over $I$ such that the induced bounded elementary embedding $V(X)\to V(^{*}X)$ is an enlargement, where $^{*}X=X^{I}/\mathcal{U}$. The idea behind the proof of the existence of a polysaturated extension is to iterate this construction.
Let $V(X)$ be a superstructure over a base set $X$, and let $\kappa=|V(X)|$ be its cardinality. For $\alpha\leq\beta\leq\kappa^{+}$, we construct superstructures $V(X^{\alpha})$ and bounded elementary embeddings $_{\alpha\beta}:V(X^{\alpha})\to V(X^{\beta})$ such that
@@
    If $\alpha\leq\kappa^{+}$, then $_{\alpha\alpha}=\text{id}_{V(X^{\alpha})}$.
@@
    If $\alpha\leq\beta\leq\gamma\leq\kappa^{+}$, then $_{\beta\gamma}\circ*_{\alpha\beta}=*_{\alpha\gamma}$.
@@
    If $\alpha\leq\beta\leq\kappa^{+}$, then $_{\alpha\beta}(X^{\alpha})=X^{\beta}$.
@@
We define $V(X^{0})=V(X)$, and $_{00}=\text{id}_{V(X)}$.
Suppose now that superstructures $V(X^{\alpha})$ and bounded elementary embeddings $_{\alpha\beta}:V(X^{\alpha})\to V(X^{\beta})$ have already been constructed for $\alpha\leq\beta<\lambda$, such that (i)-(iii) are satisfied. We now consider two cases:
Case 1: $\lambda$ is a limit ordinal. In that case construct $V(X^{\lambda})$ and $_{\alpha\lambda}$ as a limit of the $V(X^{\alpha})$ and $_{\alpha\beta}$ for $\alpha,\beta<\lambda$. We have just seen that such a construction yields bounded elementary embeddings which preserve properties (i)-(iii).
Case 2: $\lambda=\gamma+1$ is a successor ordinal. In that case, let $V(X^{\gamma})\overset{*_{\gamma\lambda}}{\to}V(X^{\lambda})$ be an enlargement, as provided by Theorem 6.7. Observe that $X^{\lambda}={}^{*_{\gamma\lambda}}X^{\gamma}$ by construction. For $\alpha<\gamma$, define $_{\alpha\lambda}=*_{\gamma\lambda}\circ*_{\alpha\gamma}$, and define $_{\lambda\lambda}=\text{id}_{V(X^{\lambda})}$. It is then straightforward to show that that properties (i)-(iii) are satisfied.
@@
For $(a,\alpha)\in P_{0}$, we define $a,\alpha]:=\{(b,\beta):(b,\beta)\sim(a,\alpha)\}$ and for $(a,\alpha)\in P-P_{0}$ we put $a,\alpha]:=\{[b,\beta]:(b,\beta)E(a,\alpha)$. Then we define
@@
$a,\alpha]=[*_{\alpha\beta}(a),\beta]\quad\text{ for }\alpha\leq\beta<\kappa^{+}.$
@@
$_{\alpha\kappa^{+}}:V(X^{\alpha})\to V(X^{\kappa^{+}}):a\mapsto[a,\alpha].$
Let $\mathcal{A}\subseteq W_{n}$ be a family of cardinality $\kappa$ with the f.i.p. We will show that $\bigcap\mathcal{A}\neq\varnothing$. Note that each $A\in\mathcal{A}$ is of the form $a,\alpha]$ for some $a$ in some $V_{n}(X^{\alpha})$. Suppose that
[…]
Let $\beta:=\sup_{\xi<\kappa}\alpha_{\xi}$. Since $\kappa^{+}$ is a regular cardinal and each $\alpha_{\xi}<\kappa^{+}$, we have that $\beta<\kappa^{+}$. Define $a_{\xi}^{\prime}=*_{\alpha_{\xi}\beta}(a_{\xi})\in V_{n}(X^{\beta})$, so that $A_{\xi}=[a_{\xi},\alpha_{\xi}]=[a^{\prime}_{\xi},\beta]$, and let $\mathcal{A}^{\prime}:=\{a^{\prime}_{\xi}:\xi<\kappa\}$ Observe that $_{\beta\kappa^{+}}(a^{\prime}_{\xi})=[a^{\prime}_{\xi},\beta]=A_{\xi}$. As $\mathcal{A}$ has the f.i.p. and $_{\beta\kappa^{+}}$ is a bounded elementary embedding, it follows easily that $\mathcal{A}^{\prime}\subseteq V_{n}(X^{\beta})$ has the f.i.p. as well. As $_{\beta(\beta+1)}:V(X^{\beta})\to V(X^{\beta+1})$ is an enlargement, it follows that
@@
Now if $x\in\bigcap*_{\alpha(\alpha+1}[\mathcal{A}^{\prime}]$, then $x\in*_{\beta(\beta+1)}a_{\xi}^{\prime}$ for all $\xi<\kappa$. It follows that $_{(\beta+1)\kappa^{+}}(x)\in*_{(\beta+1)\kappa^{+}}\circ*_{\beta(\beta+1)}(a_{\xi}^{\prime})=*_{\beta\kappa^{+}}(a^{\prime}_{\xi})=A_{\xi}$ for all $\xi<\kappa$, and hence that $_{(\beta+1)\kappa^{+}}(x)\in\bigcap\mathcal{A}$. Thus $\bigcap\mathcal{A}\neq\varnothing$ whenever $\mathcal{A}\subseteq W_{n}$ is a family of $\leq\kappa$-many sets with the f.i.p.
Observe that $X^{\kappa^{+}}=*_{\alpha\kappa^{+}}(X_{\alpha})$ for all $\alpha<\kappa^{+}$. We now show that each $_{\alpha\kappa^{+}}:V(X^{\alpha})\to V(X^{\kappa^{+}})$ is $\kappa^{+}$-saturated.
@@
Since $\kappa=|V(X)|$, it follows that if we define $_{0\kappa^{+}}$ and $^{*}X:=X^{\kappa^{+}}$, then the map $V(X)\to V({}^{*}X)$ is polysaturated.